# the eXTP mission

### v1

@@ Abstract
In this paper we present the current status of the enhanced X-ray Timing and Polarimetry mission, which has been fully approved for launch in 2030. eXTP is a space science mission designed to study fundamental physics under extreme conditions of matter density, gravity, and magnetism. The mission aims at determining the equation of state of matter at supra-nuclear density, measuring effects of QED, and understanding the dynamics of matter in strong-field gravity. In addition to investigating fundamental physics, the eXTP mission is poised to become a leading observatory for time-domain and multi-messenger astronomy in the 2030s, as well as providing observations of unprecedented quality on a variety of galactic and extragalactic objects. After briefly introducing the history and a summary of the scientific objectives of the eXTP mission, this paper presents a comprehensive overview of: 1) the cutting-edge technology, technical specifications, and anticipated performance of the mission’s scientific instruments; 2) the full mission profile, encompassing spacecraft design, operational capabilities, and ground segment infrastructure.
@@ 1 Introduction
The enhanced X–ray Timing and Polarimetry mission – eXTP is a scientific space mission designed to study the state of matter under extreme conditions of density, gravity and magnetism [1, 2, 3]. Primary goals are the determination of the equation of state of matter at supra-nuclear density, the measurement of QED effects in the radiation emerging from highly magnetized stars, and the study of matter dynamics in the strong-field regime of gravity. The matter inside neutron stars (NSs), the space-time close to Black Holes (BHs), and the extremely magnetized vacuum close to magnetars are among the uncharted territories of fundamental physics. The eXTP mission will revolutionize these areas of fundamental research by high precision X-ray measurements of NSs across the magnetic field scale and BHs across the mass scale. In addition to investigating fundamental physics, eXTP will be a very powerful observatory for the time-domain and multi-messenger astronomy in the 2030’s, as well as providing observations of unprecedented quality on a variety of galactic and extragalactic objects. The eXTP science case is described in the other five white papers included in this special issue [4, 5, 6, 7, 8]. They address the four main scientific objectives of the mission and the observatory science. These are updated substantially from the white papers published in 2019 [9, 10, 11, 12], to account for the rapid developments of the field over the past several years and the new mission profile described in the present paper.
The eXTP mission is an enhanced version of the Chinese X-ray Timing and Polarimetry mission [13], which in 2011 was selected and funded for a Phase 0/A study as one of the background concept missions in the Strategic Priority Space Science Program of the Chinese Academy of Sciences (CAS). Also in 2011, the Large Observatory for Timing (LOFT) mission concept [14, 15] was selected for an assessment study in the context of the ESA’s Announcement of Opportunity for the third of the medium size missions (M3) foreseen in the framework of the Agency’s Cosmic Vision program. The LOFT study was carried out in 2011-2014 by a consortium of European institutes. Eventually, the exoplanetary mission PLATO, considered more well timed, was selected as the ESA Cosmic Vision M3 mission. Following this, in 2015, the European LOFT consortium and the Chinese team merged the concepts of the LOFT and XTP missions, thus starting the eXTP project.
@@
According to the adopted mission profile, the Large Area Detector (LAD) and the Wide Field Monitor (WFM) are not in the baseline. In addition, the number of telescopes of the Spectroscopy Focusing Array (SFA) and Polarimetry Focusing Array (PFA) has been reduced from 9 to 6 and from 4 to 3, respectively, to facilitate deployment using a smaller, more cost-effective launch vehicle; the focal plane camera of one of the six SFA telescopes is replaced by a pnCCD camera, while the other five still use the original silicon drift detectors (SDDs). The baseline orbit has been modified from a near-equatorial low-Earth orbit (LEO) to a highly elliptical orbit (HEO) to enhance observational efficiency. However, a LEO configuration remains a viable alternative. A Wide Field and Wideband Camera (W2C) is considered as a secondary instrument to partially mitigate the loss of WFM capabilities. The eXTP consortium is led by the Institute of High Energy Physics (IHEP) of the CAS. In total, more than 200 scientists in more than 80 institutions from 17 countries are members of the eXTP instrument teams and science working groups. In this paper, we present: (1) a summary of the scientific objectives of the eXTP mission in Section 2; (2) the technological implementation and expected performance characteristics of the scientific payload instruments in Section 3; and (3) the key operational elements and functional architecture of the mission in Section 4.
@@ 2 Scientific objectives of the eXTP mission
The scientific objectives of eXTP can be divided into five themes: 1) to investigate the equation of state (EoS) of ultra-dense matter in NSs, presented in detail in White Paper 2 (WP2, [4]); 2) to understand black holes by probing strong gravity regions through synergistic X-ray spectral, timing, and polarimetric observations, presented in details in White Paper 3 (WP3, [5]); 3) to resolve some long-standing questions about QED effects and radiative processes in the strong magnetic fields, presented in details in White Paper 4 (WP4, [6]); 4) to play a key role in the era of time-domain and multi-messenger astronomy, presented in details in White Paper 5 (WP5, [7]); and 5) to provide observations of unprecedented quality on a variety of galactic and extragalactic objects, as a very powerful observatory for astrophysics, presented in details in White Paper 6 (WP6, [8]).
@@
WP2 focuses on how to determine the relation between the mass and radius of NSs with the eXTP observations, which are the key elements to studying EoS. Several methods of mass and radius measurements are proposed by different phenomena including: (1) mass and radius measurements by pulse profile modeling of rotation powered milli-second pulsars (MSPs); (2) mass and radius measurements by pulse profile modeling of accreting powered MSPs together with the polarized information; (3) mass and radius constraints by thermonuclear burst cooling tails, accretion disk dynamics and cooling curve of MSPs; (4) the inner structure of NS revealed by glitches; and (5) EoS constraints with other phenomena and methods.
@@
WP3 studies on how to understand black holes and the physics in strong gravity with eXTP observations. Leveraging its advanced instruments, eXTP will enable unprecedented precision in measuring black hole parameters and testing fundamental physics. The scientific objectives include: (1) spin and mass measurements using relativistic reflection spectroscopy (e.g., Fe line profiles), thermal continuum fitting or polarization data; (2) measuring spins and masses of supermassive black holes (SMBHs) with QPOs, X-ray reverberation in super-Eddington accretion flows, and track orbital hotspots; (3) testing general relativity by analyzing possible deviations from Kerr metrics; and (4) accretion physics, such as disk-corona dynamics and coronal geometry, with time-resolved spectral-timing techniques and polarization data.
@@
WP6 showcases that eXTP, as a very powerful observatory for astrophysics, will provide observations of unprecedented quality on a variety of galactic and extragalactic objects. The observatory science objectives may include: characterizing stellar coronae and transient phenomena such as flares and coronal mass ejections, probing magnetic turbulence and particle acceleration in supernova remnants and pulsar wind nebulae, investigating accretion physics across diverse systems (CVs, XRBs, AGNs), exploring super-Eddington accretion in ultraluminous X-ray sources (ULXs), and advancing X-ray pulsar-based timing techniques.
@@ 3.1 Spectroscopy Focusing Array - SFA
The Spectroscopic Focusing Array (SFA) comprises six Wolter-I grazing-incidence X-ray telescopes designed for spectral, timing, and imaging observations in the 0.5 to 10 keV energy range. Each telescope assembly includes the mirror module, the electron deflector, the filter wheel, and the focal plane camera. Five of the focal plane cameras are equipped with silicon drift detectors (SDDs), and their corresponding telescopes are designated SFA-T (where ’T’ denotes ’Timing’), reflecting the SDDs’ excellent timing resolution. A schematic of the SFA-T telescope design is presented in Figure 2. The sixth telescope features a pnCCD focal plane camera, analogous to those employed in eROSITA [16] and Einstein Probe [17], providing enhanced point-source sensitivity (via reduced background) and superior imaging performance. This unit is designated SFA-I (where ‘I’ signifies ‘Imaging’). The focal plane detector modules for both the SFA-T and SFA-I detectors will be developed and contributed to the eXTP project by the Max Planck Institute for Extraterrestrial Physics (MPE). The sensor chips are fabricated by the Semiconductor Laboratory of the Max Planck Society (MPG-HLL). The total effective area of SFA is expected to be larger than cm2 at 1.5 keV and cm2 at 6 keV, and the field of view (FoV) is 18 arcmin. Each SFA-T telescope includes a 19-cell SDD array, whose energy resolution is better than 180 eV at 6 keV. The time resolution is 10 s, and the dead time is expected to be less than 5% for a source flux of approximately 1 Crab. The angular resolution must be less than 1 arcmin (half-power diameter, HPD). To meet the requirements, the working temperature of the mirror modules must be stable at 20 2 ∘C. An electron deflector is installed at the bottom of the mirror module to reduce the particle background.
@@ 3.1.1 Optics
Since they are based on very similar concepts, in this section, we discuss the optics of both the SFA and the PFA together. Nine X-ray grazing-incidence Wolter I (parabola + hyperbola) optics modules will be implemented onboard eXTP; the mirror technique implemented is nickel replication, which has already been successfully used for high-throughput X-ray telescopes with good angular resolution, such as for the BeppoSAX [18], XMM-Newton [19], Swift [20], eROSITA [16] and ART-X telescopes onboard the Spectrum X-GAMMA Mission [21], the IXPE [22, 23] polarimetric mission, and the FXT telescope aboard the Einstein Probe mission [17]. The SFA includes six optic modules, while three optic modules are used for the PFA. The expected collecting area of each telescope is cm2 (SFA) and cm2 (PFA) in 3 keV, and cm2 at 6 keV. The requirement for the angular resolution of the PFA optics is arcsec (HPD), more demanding than the angular resolution of 1 arcmin (HPD) required for the SFA optics. For both types of telescope, the focal length reference value is 525 cm. The maximum diameter of the mirror shells, allowed by the envelope for a single mirror less than 600 mm in diameter, is about 510 mm. In Table 1 we list the main parameters of the SFA and PFA optics.
@@
| Angular resolution SFA | (HPD) |
| Angular resolution PFA | (HPD) |
@@ 3.1.2 SFA-T Detectors
The focal plane detector for each SFA-T telescope consists of an SDD sensor arranged in 19 hexagonal cells, with readout provided by custom ASICs, three of which are needed for each 19-cell SDD (Altmann et al. 2024). This configuration meets the aforementioned requirements (see Fig. 5). Compared with CCDs, SDDs allow a much faster readout. The detector geometry ensures that the vast majority of source photons are focused onto the inner 7 cells of the array, whereas the cells of the outermost ring are used to accurately determine the background. Thus, the side length of a hexagonal cell has been determined to be 3.2 mm matching the expected angular resolution of the optics. The sensor has a 450 m sensitive thickness, which delivers excellent quantum efficiency throughout the SFA energy range. The signal processing chain of each SDD cell includes a charge-sensitive preamplifier, a fast shaper, and a slow shaper with sample-and-hold circuit. An FPGA-based back-end electronics (BEE) module is implemented to control the front-end electronics (FEE) and read the digitized data from the ADCs.
A filter wheel is located above the detector, with four positions corresponding to 4 operational modes: (1) Open filter, used in ground tests and weak sources observations; (2) Calibration. An Fe-55 radioactive source is used for in orbit calibration; (3) Optical blocking filter. This filter, which consists of a 400 nm polyimide and a 200 nm aluminum film, is used to block visible and UV light. It also prevents contamination of the detector; (4) Closed filter. A metal shutter is used to measure the internal background and prevents the detector from being damaged during launch and extreme solar events. To compensate for particle damage to the SDDs, especially due to protons, a temperature of C is expected. during the operation of the detectors. Temperature stability of 0.5 ∘C is also required.
@@
| Energy Resolution (FWHM) | 180 eV@1.5 keV |

### v2

@@ Abstract
In this paper we present the current status of the enhanced X-ray Timing and Polarimetry mission, which has been fully approved for launch in 2030. eXTP is a space science mission designed to study fundamental physics under extreme conditions of matter density, gravity, and magnetism. The mission aims at determining the equation of state of matter at supra-nuclear density, measuring the effects of quantum electro-dynamics, and understanding the dynamics of matter in strong-field gravity. In addition to investigating fundamental physics, the eXTP mission is poised to become a leading observatory for time-domain and multi-messenger astronomy in the 2030’s, as well as providing observations of unprecedented quality on a variety of galactic and extragalactic objects. After briefly introducing the history and a summary of the scientific objectives of the eXTP mission, this paper presents a comprehensive overview of: 1) the cutting-edge technology, technical specifications, and anticipated performance of the mission’s scientific instruments; 2) the full mission profile, encompassing spacecraft design, operational capabilities, and ground segment infrastructure.
@@ 1 Introduction
The enhanced X–ray Timing and Polarimetry mission – eXTP is a scientific space mission designed to study the state of matter under extreme conditions of density, gravity and magnetism [1, 2, 3]. Primary goals are the determination of the equation of state of matter at supra-nuclear density, the measurement of Quantum Electro-Dynamics (QED) effects in the radiation emerging from highly magnetized stars, and the study of matter dynamics in the strong-field regime of gravity. The matter inside neutron stars (NSs), the space-time close to Black Holes (BHs), and the extremely magnetized vacuum close to magnetars are among the uncharted territories of fundamental physics. The eXTP mission will revolutionize these areas of fundamental research by high precision X-ray measurements of NSs across the magnetic field scale and BHs across the mass scale. In addition to investigating fundamental physics, eXTP will be a very powerful observatory for the time-domain and multi-messenger astronomy in the 2030’s, as well as providing observations of unprecedented quality on a variety of galactic and extragalactic objects. The eXTP science case is described in the other five white papers included in this special issue [4, 5, 6, 7, 8]. They address the four main scientific objectives of the mission and the observatory science. These are updated substantially from the white papers published in 2019 [9, 10, 11, 12], to account for the rapid developments of the field over the past several years and the new mission profile described in the present paper.
The eXTP mission is an enhanced version of the Chinese X-ray Timing and Polarimetry mission (XTP) [13], which in 2011 was selected and funded for a Phase 0/A study as one of the background concept missions in the Strategic Priority Space Science Program of the Chinese Academy of Sciences (CAS). Also in 2011, the Large Observatory for Timing (LOFT) mission concept [14, 15] was selected for an assessment study in the context of the ESA’s Announcement of Opportunity for the third of the medium size missions (M3) foreseen in the framework of the Agency’s Cosmic Vision program. The LOFT study was carried out in 2011-2014 by a consortium of European institutes. Eventually, the exoplanetary mission PLATO [16], considered more well timed, was selected as the ESA Cosmic Vision M3 mission. Following this, in 2015, the European LOFT consortium and the Chinese team merged the concepts of the LOFT and XTP missions, thus starting the eXTP project.
@@
According to the adopted mission profile, the Large Area Detector (LAD, a large array of collimated detectors made of silicon draft detectors (SDDs)) and the Wide Field Monitor (WFM, six modules of 1.5-D coded mask imager with SDDs as the pixelated detector planes) are not in the baseline. Both LAD and WFM are the main instruments of the LOFT mission [14, 15] and also the proposed European contributions to the previous baseline of the eXTP mission [2, 3]. In addition, the number of telescopes of the Spectroscopy Focusing Array (SFA) and Polarimetry Focusing Array (PFA) has been reduced from 9 to 6 and from 4 to 3, respectively, to facilitate deployment using a smaller, more cost-effective launch vehicle; the focal plane camera of one of the six SFA telescopes is replaced by a pnCCD (Charge Coupled Device) camera, while the other five still use the original SDDs. The baseline orbit has been modified from a near-equatorial low-Earth orbit (LEO) to a high elliptical orbit (HEO) to enhance observational efficiency. However, a LEO configuration remains a viable alternative. A Wide Field and Wideband Camera (W2C) is considered as a secondary instrument to partially mitigate the loss of WFM capabilities. The key performance parameters of the eXTP satellite’s current design are summarized in Table 1, with comparative metrics from other flagship X-ray observatories listed for contextual benchmarking.
The eXTP consortium is led by the Institute of High Energy Physics (IHEP) of the CAS. In total, more than 200 scientists in more than 80 institutions from 17 countries are members of the eXTP instrument teams and science working groups. In this paper, we present: (1) a summary of the scientific objectives of the eXTP mission in Section 2; (2) the technological implementation and expected performance characteristics of the scientific payload instruments in Section 3; and (3) the key operational elements and functional architecture of the mission in Section 4.
@@ 2 Scientific objectives of the eXTP mission
| Instrument | |||||||
| (cm2) | (keV) | (MDP) | (cts/s/cm2/keV) | (cts/s) | |||
| eXTP/SFA-T | 2750 @ 1.5 keV | 0.5–10 | s | No | No | ||
| 1670 @ 6 keV | |||||||
| eXTP/SFA-I | 550 @ 1.5 keV | 0.5–10 | 50 ms (full) | No | 3 (FF) | ||
| @ 6 keV | 240 s (timing) | ||||||
| eXTP/PFA | 250 @ 3 keV | 2–8 | s | 2% (100 ks) | No | ||
| eXTP/W2C | 160 @ 60 keV | 30–600 | ms | No | No | ||
| RXTE/PCA[17] | 6500 @ 6 keV | 2–60 | s | No | No | ||
| @ 3 keV | |||||||
| Chandra (ACIS-I)[18] | @ 1.5 keV | 0.2–10 | 3.2s | No | 0.03 | ||
| XMM-Newton (EPIC-PN)[19] | @ 1.5 keV | 0.15–12 | 30 s (timing) | No | 8 | ||
| @ 6 keV | |||||||
| NICER[20] | @ 1.5 keV | 0.2–12 | s | No | (non-imaging) | No | |
| @ 6 keV | |||||||
| IXPE[21] | @ 3 keV | 2–8 | s | ( 100 ks) | No | ||
| EP-FXT[22] | @ 1.25 keV | 0.5–10 | s | - | 5 (FF) | ||
| NewAthena/X-IFU[23] | @ 1 keV | 0.2–12 | 10 s | No | - | ||
| NewAthena/WFI[23] | @ 1 keV | 0.2–15 | 80 s (fast) | No | - | ||
| 160 s (normal) |
@@
The scientific objectives of eXTP can be divided into five themes: 1) to investigate the equation of state (EoS) of ultra-dense matter in NSs, presented in detail in White Paper 2 (WP2, [4]); 2) to understand black holes by probing strong gravity regions through synergistic X-ray spectral, timing, and polarimetric observations, presented in details in White Paper 3 (WP3, [5]); 3) to resolve some long-standing questions about QED effects and radiative processes in the strong magnetic fields, presented in details in White Paper 4 (WP4, [6]); 4) to play a key role in the era of time-domain and multi-messenger astronomy, presented in details in White Paper 5 (WP5, [7]); and 5) to provide observations of unprecedented quality on a variety of galactic and extragalactic objects, as a very powerful observatory for astrophysics, presented in details in White Paper 6 (WP6, [8]). Scientific goals in themes 1-4 are the core objectives of the current eXTP mission, expanded from the original themes 1-3 [2], which have driven the eXTP baseline mission design and its evolution.
@@
WP2 focuses on how to determine the relation between the mass and radius of NSs with the eXTP observations, which are the key elements to studying EoS. Several methods of mass and radius measurements are proposed by different phenomena including: (1) mass and radius measurements by pulse profile modeling of rotation powered milli-second pulsars (MSPs); (2) mass and radius measurements by pulse profile modeling of accretion-powered MSPs together with the polarized information; (3) mass and radius constraints by thermonuclear burst cooling tails, accretion disk dynamics and cooling curve of MSPs; (4) the inner structure of NS revealed by glitches; and (5) EoS constraints with other phenomena and methods.
@@
WP3 studies on how to understand black holes and the physics in strong gravity with eXTP observations. Leveraging its advanced instruments, eXTP will enable unprecedented precision in measuring black hole parameters and testing fundamental physics. The scientific objectives include: (1) spin and mass measurements using relativistic reflection spectroscopy (e.g., Fe line profiles), thermal continuum fitting or polarization data; (2) measuring spins and masses of supermassive black holes (SMBHs) with quasi-periodic oscillations (QPOs), X-ray reverberation in super-Eddington accretion flows, and track orbital hotspots; (3) testing general relativity by analyzing possible deviations from Kerr metrics; and (4) accretion physics, such as disk-corona dynamics and coronal geometry, with time-resolved spectral-timing techniques and polarization data.
@@
WP6 showcases that eXTP, as a very powerful observatory for astrophysics, will provide observations of unprecedented quality on a variety of galactic and extragalactic objects. The observatory science objectives may include: characterizing stellar coronae and transient phenomena such as flares and coronal mass ejections, probing magnetic turbulence and particle acceleration in supernova remnants and pulsar wind nebulae, investigating accretion physics across diverse systems (e.g., Cataclysmic Variables (CVs), X-ray binaries (XRBs), Active Galactic Nuclei (AGNs)), exploring super-Eddington accretion in ultraluminous X-ray sources (ULXs), and advancing X-ray pulsar-based timing techniques.
@@ 3.1 Spectroscopy Focusing Array - SFA
The Spectroscopic Focusing Array (SFA) comprises six Wolter-I grazing-incidence X-ray telescopes designed for spectral, timing, and imaging observations in the 0.5 to 10 keV energy range. Each telescope assembly includes the mirror module, the electron deflector, the filter wheel, and the focal plane camera. Five of the focal plane cameras are equipped with SDDs, and their corresponding telescopes are designated SFA-T (where ’T’ denotes ’Timing’), reflecting the SDDs’ excellent timing resolution. A schematic of the SFA-T telescope design is presented in Figure 2. The sixth telescope features a pnCCD focal plane camera, analogous to those employed in eROSITA [24] and Einstein Probe [22], providing enhanced point-source sensitivity (via reduced background) and superior imaging performance. This unit is designated SFA-I (where ‘I’ signifies ‘Imaging’). The focal plane detector modules for both the SFA-T and SFA-I detectors will be developed and contributed to the eXTP project by the Max Planck Institute for Extraterrestrial Physics (MPE). The sensor chips are fabricated by the Semiconductor Laboratory of the Max Planck Society (MPG-HLL). The total effective area of SFA is expected to be larger than cm2 at 1.5 keV and cm2 at 6 keV, and the field of view (FoV) is 18 arcmin. Each SFA-T telescope includes a 19-cell SDD array, whose energy resolution is better than 180 eV at 6 keV. The time resolution is 10 s, and the dead time is expected to be less than 5% for a source flux of approximately 1 Crab. The angular resolution must be less than 1 arcmin (half-power diameter, HPD). To meet the requirements, the working temperature of the mirror modules must be stable at 20 2 ∘C. An electron deflector is installed at the bottom of the mirror module to reduce the particle background.
@@ 3.1.1 Optics
Since they are based on very similar concepts, in this section, we discuss the optics of both the SFA and the PFA together. Nine X-ray grazing-incidence Wolter I (parabola + hyperbola) optics modules will be implemented onboard eXTP; the mirror technique implemented is nickel replication, which has already been successfully used for high-throughput X-ray telescopes with good angular resolution, such as for the BeppoSAX [25], XMM-Newton [19], Swift [26], eROSITA [24] and ART-XC telescopes onboard the Spektr-Roentgen-Gamma Mission [27], the IXPE [28, 21] polarimetric mission, and the FXT telescope aboard the Einstein Probe mission [22]. The SFA includes six optic modules, while three optic modules are used for the PFA. The expected collecting area of each telescope is cm2 (SFA) and cm2 (PFA) at 3 keV, and cm2 at 6 keV. The requirement for the angular resolution of the PFA optics is arcsec (HPD)@3 keV, more demanding than the angular resolution of 1 arcmin (HPD)@1.5 keV required for the SFA optics. For both types of telescope, the focal length reference value is 525 cm. The maximum diameter of the mirror shells is about 510 mm, constrained by the envelope of less than 600 mm in diameter for a single mirror. In Table 2 we list the main parameters of the SFA and PFA optics.
@@
| Angular resolution SFA | (HPD)@1.5 keV |
| Angular resolution PFA | (HPD)@3 keV |
@@ 3.1.2 SFA-T Detectors
The focal plane detector for each SFA-T telescope consists of an SDD sensor arranged in 19 hexagonal cells, with readout provided by custom ASICs, three of which are needed for each 19-cell SDD [29].
This configuration meets the aforementioned requirements (see Fig. 5). Compared with CCDs, SDDs allow a much faster readout. The detector geometry ensures that the vast majority of source photons are focused onto the inner 7 cells of the array, whereas the cells of the outermost ring are used to accurately determine the background. Thus, the side length of a hexagonal cell has been determined to be 3.2 mm given the expected angular resolution of the optics, a trade-off between the number of pixels and some sampling of the mirror’s response.. The sensor has a 450 m sensitive thickness, which delivers excellent quantum efficiency throughout the SFA energy range. The signal processing chain of each SDD cell includes a charge-sensitive preamplifier, a fast shaper, and a slow shaper with sample-and-hold circuit. A Field Programmable Gate Array (FPGA) based back-end electronics (BEE) module is implemented to control the front-end electronics (FEE) and read the digitized data from the analog-to-digital converters (ADCs).
A filter wheel is located above the detector, with four positions corresponding to 4 operational modes: (1) Open filter, used in ground tests and weak sources observations; (2) Calibration. An Fe-55 radioactive source is used for in orbit calibration; (3) Optical blocking filter. This filter, which consists of a 400 nm polyimide and a 200 nm aluminum film, is used to block visible and ultra-violet (UV) light. It also prevents contamination of the detector; (4) Closed filter. A metal shutter is used to measure the internal background and prevents the detector from being damaged during launch and extreme solar events. To compensate for particle damage to the SDDs, especially due to protons, a temperature of C is expected. during the operation of the detectors. Temperature stability of 0.5 ∘C is also required.
@@
| Energy Resolution (FWHM) | 180 eV@6 keV |